# IR Bounds on Theories with Spontaneously-Broken Lorentz Symmetry

### v2

@@ 5 Discussion
In this work we have addressed the question of what is the IR characterization of analyticity/positivity bounds in theories with spontaneously broken Lorentz symmetry. We have derived bounds in terms of the ray velocity of low energy modes in the EFT, capturing exactly the bounds derived from analyticity. We studied the context of an EFT for a conformal superfluid in $D=3$, presented in Sec. 2, studying an IR model in which the $U(1)$ symmetry group is gauged and the superfluid is coupled to a gauge vector field. In this model the EFT has a mass gap, meaning that the ray velocity goes to zero for vanishing momenta. In Sec. 4 we posed IR bounds by requiring that the ray velocity doesn’t grow too fast for low momenta. In the specific, we have imposed that the ray velocity does not grow beyond the speed of gapless excitations, $c_{s}^{2}=1/2$, for momenta below the mass gap, see Eq. (54).
@@
Technically, the similarity between our bound and that of Creminelli:2022onn can be traced back to the close connection between correlators of conserved currents and propagators of the low energy modes, which we discussed in Sec. 4.2.
@@
In our approach, we consider the vector field in the IR theory to be dynamical. Conservatively, this EFT can be thought of as an auxiliary IR system from which to extract the information contained in analyticity bounds. It would be interesting to investigate whether this system can also be considered physical, as a stable low energy phase of a conformally invariant UV theory. (See e.g. Kaplan:2023fbl ; Seiberg:2024wgj ; Berezhiani:2024boz for recent progress on gauge theories at finite charge density.) In this regard, it would be non-trivial to show that one can have a phase with vanishing electric field, even using a background charge density. Another interesting point is that the EFT we study only has gapped modes. It will be useful to explain this either in terms of analogy with plasma physics, or in terms of the algebra of broken generators, see e.g. Nicolis:2012vf .
@@
In conclusion, we have learned that the requirement that modes below the mass gap treshold should be slower than the gapless modes leads to meaningful bounds on the IR coefficients. This reproduces the known analyticity bounds. Our work moves in the direction of developing a simple way to infer the UV properties of nature from IR quantities, even when the systems observed cannot be described within the most conservative assumptions.
@@
As a direct extension of our work one might consider bounds derived for conformal superfluids in Ref. Creminelli:2022onn , for other higher-order operators that we have not studied here. The analyticity bounds for these operators were derived studying two-point correlators of the stress-energy tensor. On the IR side, learning from our results one may expect that the coupling to gravity will be important in this case. Considering the propagation of modes which mix the conformal scalar and gravitational excitations may provide the necessary handle to bound the coefficients of these irrelevant operators from the IR.
@@
In the future it would also be interesting to turn our attention towards other domains where these ideas might be applicable. Indeed, while the conformal superfluid EFT that we studied here is interesting in its own right, it can also be considered as a proxy for more complicated theories in which Lorentz invariance is spontaneously broken. Such is the case of cosmological theories of modified gravity, our original motivation.

### v3

@@ V Discussion
In this work we have approached the question of what is the IR characterization of analyticity/positivity bounds in theories with spontaneously broken Lorentz symmetry. We have expressed the known bounds in terms of a momentum-independent condition on the kinematics of low energy modes in the EFT, specifically through the acoustic vector $N^{\mu}$ and the associated velocity $\bf v$}. We studied this in the context of an EFT for a conformal superfluid in $D=3$, presented in Sec. II, studying an IR model in which the $U(1)$ symmetry group is gauged and the superfluid is coupled to a gauge field. In this model the EFT has both a mass gap and HD operators, both contributing to the presence of dispersivity. In Sec. IV we recovered the known analyticity bounds for this theory by requiring that the velocity $\bf v$} from the acoustic vector $N^{\mu}$ does not grow too fast for low momenta. Specifically, we have imposed that it does not grow beyond the speed of gapless excitations, $c_{s}^{2}=1/2$, for momenta below the mass gap; see Eq. (58). In doing so, we also learned that the bound is not expressible as momentum-independent condition on neither phase nor group velocities.
@@
Technically, the the fact that the bound of [27] represents a statement about the kinematics of the gapped modes of the theory can be traced back to the close connection between correlators of conserved currents and propagators of the low energy modes, which we discussed in Sec. IV.3.
@@
In our work, we considered the IR kinematics of the gauge field excitations. Conservatively, this gauge field can be thought of as an auxiliary variable that is helpful in order to study the interactions, similarly to the approach of [27]. It would be interesting to investigate whether this system can also be considered as a physical, stable, low-energy phase of a conformally invariant UV theory. [See e.g., [46, 47, 48] for recent progress on gauge theories at finite charge density.] In this regard, it would be non-trivial to show that one can have a phase with vanishing electric field, even using a background charge density. Another interesting point is that the EFT we study only has gapped modes. It will be useful to explain this either in terms of analogy with plasma physics, or in terms of the algebra of broken generators, see e.g., [49].
@@
In conclusion, we have learned that the requirement that modes below the mass gap threshold should be slower, in a specific sense, than the gapless modes, leads to meaningful bounds on the IR coefficients. Equivalently, the energy is bounded by a combination of the mass and the energy of a gapless mode with the same spatial momentum. This is the information contained within the known analyticity bounds. Complementarily, the fact that these bounds do not express sharp statements in terms of neither phase nor group velocities of the any IR modes of the theory indicates that causality has a nuanced character in this system. Our work moves in the direction of developing a simple way to characterize the UV properties of nature from IR quantities, even when the systems observed cannot be described within the most conservative assumptions.
@@
As a direct extension of our work, one might consider bounds derived for conformal superfluids in Ref. [27], for other higher-order operators that we have not studied here. The analyticity bounds for these operators were derived studying two-point correlators of the stress-energy tensor. On the IR side, learning from our results one may expect that the coupling to gravity will be important in this case. Considering the propagation of modes which mix the conformal scalar and gravitational excitations may provide the necessary handle to recover the bound on the coefficients of these irrelevant operators from the IR. We may expect this to extend to similar but more general cases in which a symmetry can be gauged, though a full rigorous derivation of this relation is beyond the scope of this paper.
The extension of these results to other systems without Lorentz invariance requires further analysis, as many open questions remain. It would be premature to state here that analyticity bounds will in general be expressible in the IR as a momentum-independent condition on the velocity defined from the acoustic vector $N^{\mu}$. However, note that for gapless theories this requirement is the same as finding a momentum-independent bound on the phase velocity, which is known to reproduce analyticity bounds in several cases. The presence of a gap here establishes a window in momenta where the bound is effectively applied, but this may be a peculiarity that does not generalize beyond the conformal superfluid in $d=3$. Still, our result confirms the expectation that consistent interactions slow down excitations with respect to the free theory, which may still be the case for more general systems. What is clear from our results is that analyticity bounds do not always take the form of a simple comparison between the speed of IR excitations and the speed of light. Instead, the acoustic metric allows to capture more nuanced properties of the IR kinematics that synthetically express analyticity bounds.
@@
In the future, it would also be interesting to turn our attention toward other domains where these ideas might be applicable. Indeed, while the conformal superfluid EFT that we studied here is interesting in its own right, it can also be considered as a proxy for more complicated theories in which Lorentz invariance is spontaneously broken. Such is the case of cosmological theories of modified gravity, our original motivation.
@@ Acknowledgements.
Appendix A An example – P(X) theory
To see that working with the geometrical description based in the acoustic metric reproduces the usual known results, let us consider the simple example of a $P(X)$ theory with Lagrangian
As this is a second-order theory without a gap, there is no dispersivity. The corresponding acoustic metric can be obtained by expanding at quadratic order in perturbations around a fixed background, $\phi=\bar{\phi}+\varphi$, and taking the Eikonal limit. It reads as
which is gapless. We proceed to compute the inverse of $Z^{\mu\nu}$, easily found to be
This allows one to write the EoM in the geometric optics approximation in terms of the acoustic vector $N^{\mu}$, obtaining
From here we can compute the velocity as defined in Eq. (37),
which coincides with the usual definition of phase velocity $\omega^{2}({\bf p})/{\bf p}^{2}$ for a background with a spatially isotropic acoustic metric. [^9] See the discussion in [40]. If we impose $N^{\mu}p_{\mu}=0$, then it follows that $\frac{{\bf N}\cdot{\bf p}}{N^{0}\omega}=1$, meaning that $\omega^{2}/{\bf p}^{2}={\bf N}^{2}/(N^{0})^{2}$, provided $\bf p$} is parallel to $\bf N$}.
Suppose that in this theory we want to avoid superluminality. Requiring the velocity (82) to be less or equal to the speed of light $c=1$ immediately implies [^10] Here we are assuming that the denominator above is always positive, which is guaranteed by the absence-of-ghosts condition.
$c_{2}\geq 0\,.$ (83)
This is the well known condition to prevent superluminality in this model [6]. Geometrically, this means the acoustic vector $N^{\mu}$ must be timelike (or null) according to the standard spacetime metric $\eta_{\mu\nu}$, i.e. lie inside (or on) the lightcone. Using Eq. (81) above
which again implies (83).
As a closing remark, in this case due to the absence of a mass term, we could have also obtained this result by instead demanding that the momentum co-vector $p_{\mu}$ lies outside (or on) the lightcone, that is, it is $\eta$-spacelike (or null). From the EoM in terms of $p_{\mu}$, we get
again leading to (83). However, in general this does not imply that the velocity in Eq. (37) agrees with $\omega^{2}({\bf p})/{\bf p}^{2}$. This is especially important when comparing it against $c_{s}^{2}\neq 1$ or when handling a theory with a mass gap.